OpenStudy (anonymous):

For the function below, state the amplitude, period, phase shift, and vertical translation of the graph for the sinusoid with respect to the parent function F(x)=sin(x). H(x)=-3sin(x- pi/4) +2 What does it mean "with respect to the parent function"?

OpenStudy (anonymous):

Basically what it is saying is what has been added to the F(x) function. The F(x) is the basic and the H(x) is the modified. I don't know if that makes sense..

OpenStudy (anonymous):

So, I wouldn't have to take in account F(x)=sin(x) when finding the amplitude, period, etc|

OpenStudy (anonymous):

You have to take in account F(x)=sin(x) because it helps you to determine the changes (amplitude, Period, etc.) that have been applied in H(x).

OpenStudy (anonymous):

So what I'm trying to find is the difference in amplitude, period, etc. between H(x) and F(x)?

OpenStudy (anonymous):

So, graphs in the form H(x) = A sin (Bx - C) +D are transformations of f(x) = sin x Do you understand this?

OpenStudy (anonymous):

Yes, I do

OpenStudy (anonymous):

Do you know which one is which in regard to amplitude, period, phase shift and vertical translation?

OpenStudy (anonymous):

Yes. A is Amplitude, B is used in finding the period, phase shift is C, and vertical shift is D, correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So, what is it that you're not understanding?

OpenStudy (anonymous):

I just don't know what it's asking for when it says find the amplitude, etc. in relation to the parent function. Am I just looking for the all of that for the function of H(x), or is there something more?

OpenStudy (anonymous):

You're just stating the transformations (amplitude, period etc.) in H(x) that have been applied to the f(x) function.

OpenStudy (anonymous):

OH. I just realized what it was asking. The "respect to the parent function" part is just stating that sin(x) is the base function, without any phase shift or amplitude changes applied, right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Hm, just to clarify, could I attach a picture of the exact question I have so I know if that's the case?

OpenStudy (anonymous):

http://gyazo.com/48615349b871c2667b0fa3cc871f6af3 That is the question above ^. Is it just asking "Identify the amplitude, period, and phase shift of the function with respect to the parent function."?

OpenStudy (anonymous):

That is what i'm getting from it.

OpenStudy (anonymous):

Alright, thank you for your help!